# ANote[fnsymbol]

### v2

@@ 3 Speech Synthesis 40M
For the non-speech datasets, we employ the zero-shot in-context generation technique proposed by CosyVoice (Du et al., 2024) to convert the text into natural-sounding speech. To enhance the diversity of voices, we randomly select a speaker from a pool of two thousand individuals for each corresponding language to synthesise the speech, ensuring that the voice aligns with the language of the text. However, many conversational and instruction datasets are not ideal for speech synthesis due to the nature of the tasks, which do not facilitate effective speech interaction. To address this issue, we enhance the quality of our synthesised speech data by implementing the following optimisation measures:
  • Length Filtering: texts exceeding a predefined length threshold are excluded from the dataset. Specifically, texts longer than 200 Chinese characters or 200 English words are removed.
@@
We further enhance the data quality with the help of a large model through the following steps:
  • Filter out questions that are not suitable for oral QA.
  • Rewrite the command in a conversational tone and generate a conversational response.
@@ 3.3 Training Strategies
Figure 4: Overview of the training stage in Nexus-O. The first stage aims to map the speech features into semantic space, the second stage aims to enable the audio instruction-following capability, and the last stage aims to enable speech generation capability.
@@
The Dataset Format Example of ChatML <|im_start|>system You are a helpful assistant.<|im_end|> <|im_start|>user <|audio_bos|>audio.wav<|audio_eos|>Please help me transcribe my current speech into text. <|im_end|> <|im_start|>assistant hat had its source away back in the woods of the old cuthbert place it was reputed to be an intricate headlong brook in its earlier course through those woods with dark secrets of pool and cascade but by the time it reached lynde’s hollow it was a quiet well conducted little stream <|im_end|>
Stage 2: audio SFT.
The second stage aims to enable the model to follow speech instructions by incorporating speech-based questions and text-based answers. In this stage, we freeze the visual and audio encoders while unfreezing the visual and audio adapters, and backbone LLM. This strategy is designed to enhance the capability of Nexus-O to respond accurately to multimodal instructions, thereby improving its performance in understanding and processing both visual and auditory information.
By keeping the core encoder parameters fixed, we ensure that the foundational representations remain stable while allowing the adapters to fine-tune and specialize in the nuances of specific tasks or instructions. This strategic adjustment aims to refine the sensitivity of Nexus-O to multimodal cues without disrupting the robustness of its underlying architecture.
@@
Different from the previous large speech model connected to the TTS module (Fu et al., 2024c; Li et al., 2024a), we exploit a decoder to generate speech end-to-end. First, we train Nexus-O on the Text-to-Speech dataset to make the model learn text-to-speech alignment. Then, we utilize the speech-to-speech dataset to train the model to learn speech understanding and speech synthesis capabilities. At this stage, only the parameters of the audio decoder are trainable.
3.4 Nexus-O-audio testbed
Our Nexus-O-audio testbed comprises audio samples collected from real-world scenarios, including video meetings, live streaming, voice-based customer service QA interactions, and real-life conversations. We apply an off-the-shelf ASR model to convert each speech into corresponding text as our testbed. More information is provided in Figure 3. In addition to the ASR task, we plan to introduce Audio QA, Audio Visual QA tasks, and tri-modal tasks, which are currently under development.
@@ 4.1 Multimodal Tasks Alignment Analysis
In this section, we evaluate the performance of Nexus-O on the following four tasks: vision understanding, audio English tasks, ASR, and S2TT. These tasks cover a wide range of modalities and capabilities, providing a comprehensive assessment of the versatility and effectiveness of MLLMs across various domains.
4.1.1 Vision-Language Evaluation
We first quantitatively evaluate Nexus-O on various vision understanding tasks, including the following benchmarks: (1) HallusionBench (Hal) (Guan et al., 2023): evaluating the hallucination and visual illusion. (2) MathVista (MathV) (Lu et al., 2023): evaluating mathematical reasoning in visual contexts. (3) OCRBench (OCR) (Fu et al., 2024d): evaluating Optical Character Recognition (OCR) capabilities on text-intensive images. (4) Video-MME (Fu et al., 2024b): evaluating video QA reasoning capabilities of MLLMs over 6 domains, diverse time range, etc. (5) MMMU (Yue et al., 2024): evaluating MLLMs on massive multi-discipline tasks demanding college-level subject knowledge and deliberate reasoning. (6) AI2D: AI2 Diagrams (AI2D) is a dataset of over 5000 grade school science diagrams. (7) MMVet (Yu et al., 2024): evaluating the integrated capabilities of MLLMs, including recognition, OCR, knowledge, language generation, spatial awareness, and math. (8) MME (Fu et al., 2024a): measuring both perception and cognition abilities on a total of 14 subtasks.
As illustrated in Table 2, we observe that Nexus-O achieves strong performance to current popular vision-language MLLMs on the MathV, OCR, and Hal benchmarks. Furthermore, when compared to Qwen2.5-VL, Nexus-O demonstrates superior performance, highlighting its competitive capabilities in vision understanding and reasoning tasks, indicating our model can effectively maintain vision-language alignment capability after auditory pretraining stage.
Table 2: Evaluation on Vision Understanding Benchmarks. The best two values are shown in bold and underlined. The blue row refers to the main competitors. In MMMU benchmark, * indicates that the measurement is developed over the validation set. We can observe that Nexus-O shows performance comparable to the leading open-source models and advanced closed-source counterparts.
@@ 4.1.1 Vision-Language Evaluation
Nexus-O Qwen2.5-VL-7B 57.0 53.2* 62.1 71.1 81.2 882.0 - 2315.5
4.1.2 Audio-Language Evaluation
Next, we evaluate the audio-language alignment on English QA, ASR, and S2TT tasks.
English QA task.
We evaluate the performance on the spoken QA benchmark: LLaMA Q. (Nachmani et al., 2024). As presented in Table 3, our model achieves top performance, higher than same-period competitor MiniCPM-o2.6-7B (highlighted in blue), demonstrating its competitive capabilities in this task.
@@ 4.1.2 Audio-Language Evaluation
Nexus-O Audio&Text 67.33
ASR task.
In the ASR task, we focus on both Mandarin (cn) and English (eng), evaluating performance on the following benchmarks: AIShell-2 (Du et al., 2018b), Librispeech (Panayotov et al., 2015b), and our real-scenario benchmark. As demonstrated in Table 4, our model achieves the best performance on the real scenario benchmarks, indicating the robustness of our model.
Table 4: Evaluation on ASR Benchmarks. Nexus-O has demonstrated strong performance in both real-life Mandarin and English ASR tasks. It outperforms specialized speech models, achieving better results in both languages. AIShell-2 is measured over all categories (Mac/iOS/Android), TC: test-clean, TO: test-other. We can observe that Nexus-O outperforms other models in real scenarios.
@@
AIShell-2 Nexus-O-audio Librispeech TC Librispeech TO Nexus-O-audio
@@
Nexus-O 3.35 11.82 3.50 5.29 12.32
Speech-to-Text translation task.
As for the S2TT, we exploit CoVoST2 (Wang et al., 2020), which provides massive multilingual S2TT datasets. As shown in Table 5, Nexus-O demonstrates superior performance compared to specialised speech LLM, Qwen2-Audio-7B-Instruct, in both translation tasks. Nevertheless, our model still falls short of the performance of its contemporary competitor, MiniCPM-o2.6-7B. However, we pre-train our model on only 8 million audio samples, which is a relatively modest dataset at the pretraining stage.
Table 5: Evaluation on Speech-to-Text Translation Benchmarks. Nexus-O demonstrated superior performance compared to Qwen2-Audio-Instruct-7B in both translation tasks, however, it still falls short of the performance of its contemporary competitor. We conjecture that this performance gap may be due to the relatively modest scale of our pretraining speech dataset.
@@
Qwen2-Audio-7B 24.40 45.20
Qwen2-Audio-Instruct-7B* 22.90 39.50
MinMo-7B * 25.95 46.68
@@
Nexus-O 23.17 40.21
@@ 4.2 Representational Alignment Analysis
In this section, we evaluate the alignment from the perspective of representation space to offer a deep evaluation beyond conventional downstream task alignment. We formalise MLLM within the framework of an unembedding-embedding architecture. In this framework, the unembedding stage is responsible for learning transformations between observations (e.g., text, vision, audio) and latent spaces through encoders, while the embedding stage captures the complex interactions among latent variables within the latent space of the hidden layers in MLLMs. Each stage serves distinct functions and yields representations with different properties (Park et al., 2024; Zhang et al., 2024d). Consequently, by focusing on each stage independently, we can have a systematical evaluation of model behaviours in representation spaces. To assess the representational alignment among vision-language modalities at each stage, we next measure the geometrical similarity among them via the kernel-alignment metrics. More information is provided in Appendix B.
Unembedding stage.
In the unembedding stage, we evaluate representational alignment using two language-vision datasets: ImageNet (Deng et al., 2009) and Wikipedia Caption (WIT) (Srinivasan et al., 2021). These datasets offer varying levels of fine granularity in language-vision alignment, enabling a comprehensive assessment of representational performance. To get the audio input for each language-vision pair, we apply the pretrained Text-to-Speech LLMs, cosyvoice-300M-SFT checkpoint, (Du et al., 2024). As for the baseline, we compare our model with MLLMs after SFT, including Qwen2.5-VL (Bai et al., 2025), LLaMA3.2-Vision (Meta., 2024), LLaVA family Liu et al. (2023b; a; 2024), Phi-3.5-Vision Abdin et al. (2024), and InternVL-8B (Chen et al., 2024c). The kernel-alignment metrics involve cka (Kornblith et al., 2019), cknna (Huh et al., 2024), svcca (Raghu et al., 2017), and $k$-nearest neighbours (knn)-based metrics: cycle knn, mutual knn, lcs knn, and edit knn. More information for those metrics can be found in (Huh et al., 2024).
As shown in Table 6, our tri-modal model incorporating audio inputs can generally outperform other models across both corpora. This demonstrates that the inclusion of the audio modality enhances representational alignment between the language and vision modalities. Additionally, Figure 5 provides a Principal Component Analysis (PCA) visualisation of the vector spaces for different modalities. The visualisation highlights that each modality exhibits distinct latent space geometries, and concatenating the vision and audio spaces results in a fusion of previously separated latent space structures.
@@
Figure 5: The unembedding space is visualized using Principal Component Analysis (PCA), with distinct colours representing clusters identified by k-means clustering. The visualisation reveals that, despite being trained within a unified model, the different modalities exhibit distinct latent space geometries.
Embedding stage.
Next, we focus on the vision-language alignment at the embedding stage. We restrict our attention to the Wikipedia-caption corpus (Srinivasan et al., 2021). We compare our model exclusively with Qwen2.5-VL-7B. For each sample, the representation is the averaged token embeddings or the last token embeddings. Table 7 illustrates the kernel-alignment scores at the last hidden layer, we can observe that incorporating audio into the vision modality improves the language-vision alignment of MLLMs, indicating the audio modality can provide complementary information to vision and language modalities.
Table 7: Kernel alignment analysis for embedding space. Some baselines (InternVL, Llava, etc.) do not support a single modality as input. Therefore, we do not evaluate them. We can observe that by incorporating audio into the vision modality, the vision-language leads to better alignment.
Baseline cycle knn mutual knn lcs knn cka cknna svcca edit knn
avg token embeds at last hidden layer
Qwen2.5-VL-7B 0.51660 0.08490 2.35839 0.20811 0.04076 0.46152 0.00201
Nexus-O 0.77929 0.13716 3.35058 0.20746 0.10547 0.47505 0.00431
vision+audio 0.79687 0.16107 3.76074 0.25915 0.12871 0.52091 0.00578
last token embed at last hidden layer
Qwen2.5-VL-7B 0.57421 0.09572 2.57714 0.21278 0.04942 0.34796 0.00226
Nexus-O 0.77343 0.10203 2.72851 0.10599 0.07068 0.39663 0.00306
vision+audio 0.90429 0.12000 3.12402 0.12787 0.08621 0.41922 0.00367
Additionally, we present a comparative analysis between Qwen2.5-VL-7B and our model across all hidden layers. As illustrated in Figure 6, the incorporation of the audio modality significantly enhances vision-language alignment, particularly in the intermediate hidden layers. Furthermore, we also provide a qualitative evaluation of language and vision alignment via PCA visualisation. In Figure 7, the observation reveals that the vision-audio fusion (depicted in the two images on the right) demonstrates a greater degree of overlap compared to the language-vision-audio separation (illustrated in the two images on the left). This suggests that the integration of audio enhances the alignment between language and vision modalities.
Figure 6: Averaged kernel-alignment score across different hidden layers, we can observe that audio modality (green bar) can result in better vision-language alignment at most hidden layers.
Figure 7: Qualitative evaluation for language-vision alignment at embedding hidden layers. We can observe that the vision-audio fusion (right two) has a larger overlap than language-vision-audio separation (left two), indicating incorporating audio leads to better language-vision alignment. we also present the PCA visualisation at all hidden layers in Figure 8 and 9.

### v3

@@ 3 Speech Synthesis 40M
We integrate both English and Chinese speech and textual datasets to generate paired speech-text data for ASR, speech synthesis (text-to-speech), Speech-to-Text Chat, and Speech-to-Speech Chat tasks. For the speech corpus (depicted at the bottom of Figure 3), we employ Whisper (Radford et al., 2022) to transcribe speech into text. For noisy samples in our in-house dataset, human annotators are engaged to ensure accurate transcription.
For the textural corpus (top in Figure 3), we employ the zero-shot in-context generation technique proposed by CosyVoice (Du et al., 2024b) to convert the text into natural-sounding speech. However, many conversational and instruction textural datasets are not ideal for speech synthesis due to the nature of the tasks, which do not facilitate effective speech interaction. To address this issue, we enhance the quality of our synthesised speech data by implementing the following pre-processing rules:
  1. 1.
    Length Filtering: texts exceeding a length threshold are excluded from the dataset. Specifically, texts longer than 200 Chinese characters or 200 English words are removed.
  2. 2.
@@
For Speech-to-Text Chat and Speech-to-Speech Chat tasks, some texts are unsuitable for a conversational context. To address this limitation, we implement the following rules:
  1. 1.
    Excluding queries that are not appropriate for the Chat task. For example, questions involving metathetical questions are filtered out.
  2. 2.
    Rewriting the question in a conversational tone and generating a conversational response via Qwen2-72B (Chu et al., 2024b).
To enhance the diversity of voices, we randomly select a speaker from a pool of two thousand individuals for each corresponding language to synthesise the speech, ensuring that the voice aligns with the language of the text. Following the text-to-speech (TTS) process, an automated verification procedure is executed using the pretrained Whisper model, which checks whether the synthesised sample can be approximately reconverted to the original source sample, thereby ensuring the quality of the data synthesis.
Finally, leveraging the synthesised audio dataset, we introduce the Nexus-audio testset, which comprises 8.1k Chinese and 2.8k English ASR samples. This testset spans a range of application domains, such as corporate meetings, live broadcasting scenarios, etc.
@@ 3.3 Training Strategies
Figure 4: Overview of the training stage in Nexus. The first stage aims to map the speech features into semantic space, the second stage aims to enable the audio instruction-following capability, and the last stage aims to enable speech generation capability.
@@
Input Format <|im_start|>system You are a helpful assistant.<|im_end|> <|im_start|>user <|audio_bos|>[audio.wav]<|audio_eos|>Please help me transcribe my current speech into text. <|im_end|> <|im_start|>assistant [text] <|im_end|>
Stage 2: audio SFT
The second stage aims to enable the model to follow speech instructions by incorporating speech-based questions and text-based answers. In this stage, we freeze the visual and audio encoders while unfreezing the visual and audio adapters, and backbone LLM. This strategy is designed to enhance the capability of Nexus to respond accurately to multimodal instructions, thereby improving its performance in understanding and processing both visual and auditory information.
By keeping the core encoder parameters fixed, we ensure that the foundational representations remain stable while allowing the adapters to fine-tune and specialize in the nuances of specific tasks or instructions. This strategic adjustment aims to refine the sensitivity of Nexus to multimodal cues without disrupting the robustness of its underlying architecture.
@@
Different from the previous large speech model connected to the TTS module (Fu et al., 2024c), we exploit a decoder to generate speech end-to-end. First, we train Nexus on the Text-to-Speech dataset to make the model learn text-to-speech alignment. Then, we utilize the speech-to-speech dataset to train the model to learn speech understanding and speech synthesis capabilities. At this stage, only the parameters of the audio decoder are trainable.
@@ 4.1 Multimodal Tasks Alignment Analysis
In this section, we evaluate the performance of Nexus on the following five tasks: vision understanding, Spoken English QA tasks, ASR, Speech-to-Text translation, and Text-to-Speech tasks. These tasks cover a wide range of modalities and capabilities, providing a comprehensive assessment of the versatility and effectiveness of MLLMs across various domains.
Vision-Language Evaluation
We conduct a comprehensive quantitative evaluation of Nexuson a range of vision understanding tasks using eight distinct benchmarks: (1) HallusionBench (Hal) (Guan et al., 2023): Evaluates the model’s susceptibility to hallucination and its handling of visual illusions. (2) MathVista (MathV) (Lu et al., 2023): Assesses mathematical reasoning within visual contexts. (3) OCRBench (OCR) (Fu et al., 2024d): Measures optical character recognition (OCR) capabilities on text-intensive images. (4) Video-MME (Fu et al., 2024b): Examines video question-answering reasoning abilities across six domains over diverse temporal ranges. (5) MMMU (Yue et al., 2024): Challenges the model with multi-disciplinary tasks that demand college-level subject knowledge and deliberate reasoning. (6) AI2D: A dataset comprising over 5,000 grade school science diagrams. (7) MMVet (Yu et al., 2024): Evaluates integrated capabilities including recognition, OCR, domain knowledge, language generation, spatial awareness, and mathematical reasoning. (8) MME (Fu et al., 2024a): Measures both perceptual and cognitive abilities across 14 subtasks.
As illustrated in Table 2, we observe that Nexus achieves strong performance to current popular vision-language MLLMs on the MathV, OCR, and Hal benchmarks. Furthermore, when compared to Qwen2.5-VL, Nexus demonstrates superior performance, highlighting its competitive capabilities in vision understanding and reasoning tasks, indicating our model and training strategy can effectively maintain vision-language alignment capability by incorporating audio modality (Finding1).
Table 2: Evaluation on Vision Understanding Benchmarks. The best two values are shown in bold and underlined. The blue row refers to the main competitors. In MMMU benchmark, * indicates that the measurement is developed over the validation set. We can observe that Nexus shows performance comparable to the leading open-source models and advanced closed-source counterparts.
@@
Nexus Qwen2.5-VL-7B 57.0 53.2* 62.1 71.1 81.2 882.0 - 2315.5
Audio-Language Evaluation
Next, we evaluate the audio-language alignment on Spoken Question-Answering (SQA), Automatic Speech Recognition (ASR), Speech-to-Text translation (S2TT), and Text-to-Speech (TTS) tasks.
SQA task
We evaluate the performance on the English spoken QA benchmark: LLaMA Q. (Nachmani et al., 2024). As presented in Table 3, our model achieves top performance, higher than same-period competitor MiniCPM-o2.6-7B (highlighted in blue), demonstrating its competitive capabilities in this task.
@@
Nexus Audio&Text 67.33
ASR task
In the ASR task, we focus on both Mandarin (CH) and English (EN), evaluating performance on the following benchmarks: AIShell-2 (Du et al., 2018b), Librispeech (Panayotov et al., 2015b), and our real-scenario benchmark. As demonstrated in Table 4, our model achieves the best performance on the real scenario ASR testset, Nexus-audio, indicating the robustness of our model in real life (Finding2).
Table 4: Evaluation on ASR Benchmarks. Nexus has demonstrated strong performance in both real-life Mandarin and English ASR tasks. It outperforms specialized speech models, achieving better results in both languages. AIShell-2 is measured over all categories (Mac/iOS/Android), TC: test-clean, TO: test-other. We can observe that Nexus outperforms other models in real scenarios.
@@
AIShell-2 Nexus-audio Librispeech TC Librispeech TO Nexus-audio
@@
Nexus 3.35 11.82 3.50 5.29 12.32
S2TT task
As for the S2TT, we exploit CoVoST2 (Wang et al., 2020), which provides massive multilingual S2TT datasets. As shown in Table 5, Nexus demonstrates superior performance compared to specialised speech LLM, Qwen2-Audio-7B-Instruct, in both translation tasks. Nevertheless, our model still falls short of the performance of its contemporary competitor, MiniCPM-o2.6-7B. However, we pre-train our model on a small audio dataset (about 30k hours) at the pre-training stage (Finding3).
Table 5: Evaluation on Speech-to-Text Translation Benchmarks. Nexus demonstrated superior performance compared to Qwen2-Audio-Instruct-7B in both translation tasks, however, it still falls short of the performance of its contemporary competitor due to the relatively modest scale of our pre-training speech dataset (about 30k hours).
@@
Qwen2-Audio-7B (Chu et al., 2024a) 24.40 45.20
Qwen2-Audio-Instruct-7B (Chu et al., 2024a)* 22.90 39.50
MinMo-7B (Chen et al., 2025a)* 25.95 46.68
@@
Nexus 23.17 40.21
TTS task
Finally, we evaluate Nexus on Seed-TTS benchmark Anastassiou et al. (2024) [^5] https://github.com/BytedanceSpeech/seed-tts-eval. As shown in Table 6, our experiments demonstrate that when employing pretrained vocoders (e.g., Fishspeech1.4 or CosyVoice2.0), Nexus achieves performance comparable to its competitors (e.g., CosyVoice1.0 or CosyVoice2.0) (Finding4).
Table 6: Evaluation on Text-to-Speech (TTS) Benchmarks. Nexus demonstrated lower performance in both Mandarin (CH) and English (EN) TTS tasks. Notice: MinMo-7B’s result is based on their manually filtered samples as mentioned in their report Chen et al. (2025a), therefore, we do not compare with it.
Model CH EN
[…]
CosyVoice1.0 (Du et al., 2024a) 8.64 3.59
CosyVoice2.0 (Du et al., 2024c) 4.61 2.43
CosyVoice2.0-SFT (Du et al., 2024c) 2.06 3.19
MinMo-7B (Chen et al., 2025a) 2.48 2.90
Nexus (Fishspeech1.4) 7.55 13.02
Nexus (CosyVoice2.0) 4.53 4.11
Thus far, our evaluations demonstrate that incorporating the audio modality enables the model to perform audio understanding and generation without compromising its vision comprehension and reasoning capabilities. In the following section, we provide an in-depth evaluation of the vision-language alignment from the perspective of representation space.
@@ 4.2 Representational Alignment Analysis
We formalise MLLM within the framework of an unembedding-embedding architecture. In this framework, the unembedding stage is responsible for learning transformations between observations (e.g., text, vision, audio) and latent spaces through encoders, while the embedding stage captures the complex interactions among latent variables within the latent space of the hidden layers in MLLMs. Each stage serves distinct functions and yields representations with different properties. Consequently, by focusing on each stage independently, we can have a systematical evaluation of model behaviours in representation spaces.
To assess the representational alignment among vision-language modalities at each stage, we first extract embeddings for each sample’s modalities. Our evaluation focuses on two aspects: (1) whether our model inherently promotes superior vision-language alignment, and (2) whether incorporating the audio modality into the vision modality further enhances this alignment.
Quantitative evaluation of vision-language alignment is performed using kernel-alignment metrics. The kernel of a space encapsulates its underlying geometrical structure (i.e., distance metric), and the similarity between the kernels of different spaces serves as an indicator of the alignment between these spaces—in our case, across different modalities. Those metrics involve cka (Kornblith et al., 2019), cknna (Huh et al., 2024), svcca (Raghu et al., 2017), and $k$-nearest neighbours (knn)-based metrics: cycle knn, mutual knn, lcs knn, and edit knn. More information for those metrics can be found in (Huh et al., 2024).
Unembedding stage
In the unembedding stage, we evaluate representational alignment using ImageNet (Deng et al., 2009), which offers a fine granularity in language-vision alignment, enabling a detailed assessment of representational performance. To get the audio input for each language-vision pair, we apply the pretrained Text-to-Speech LLMs, cosyvoice-300M-SFT checkpoint (Du et al., 2024b). As for the baseline, we compare our model with MLLMs after SFT, including Qwen2.5-VL (Bai et al., 2025), LLaMA3.2-Vision (Meta., 2024), LLaVA family Liu et al. (2023), Phi-3.5-Vision Abdin et al. (2024), and InternVL-8B (Chen et al., 2024).
As shown in Table 7, we can observe that (1) our model without audio inputs can lead to better vision-language alignment than the backbone model, Qwen2.5-VL-7B, on 5 out of 7 metrics , and (2) by incorporating audio modalities, our model outperforms others on 4 out of 7 metrics (Finding5). These results indirectly demonstrate that our modular formalism design concept can yield an efficient tri-modal alignment at the encoding stage.
@@
Figure 5: Averaged kernel-alignment score across different hidden layers. vision+audio: both visual and auditory modalities are concurrently fed into the model. The final fused representation is the last token in the sequence as the comprehensive summary of the integrated modalities. we can observe that incorporating audio modality (green bar) can result in better vision-language alignment at most hidden layers.
Embedding stage
Next, we focus on the vision-language alignment at the embedding stage. We restrict our attention to the Wikipedia-caption corpus (Srinivasan et al., 2021). We present a comparative analysis between Qwen2.5-VL-7B and our model across all hidden layers. We feed both image and its audio into the model and consider the last token embedding as the final fused representation. As illustrated in Figure 5, the incorporation of the audio modality significantly enhances vision-language alignment, particularly in the intermediate hidden layers (Finding6). This finding underscores the efficacy of our training strategy, whereby pretraining exclusively on audio data yields a robust tri-modal alignment.